1. The neutralization reaction between HCl and Ca(OH)2 produces CaCl2 and water. If 0.75g of Ca(OH)2 is used in the reaction, determine the following:
a) Number of moles of HCl needed to completely react with Ca(OH)2¬
b) Mass of CaCl2 formed in the reaction.
1
Expert's answer
2020-09-21T06:53:04-0400

Solution.

"m(Ca(OH)_2)=0.75g;"

"2HCl+Ca(OH)_2 \to CaCl_2+2H_2O;"

"a) M(Ca(OH)_2)=74g\/mol;"

"\nu=\dfrac{m}{M}; \nu(Ca(OH)_2)=\dfrac{0.75g}{74g/mol}=0.01mol;"

"x mol" "0.01mol"

"2HCl+Ca(OH)_2 \to CaCl_2+2H_2O;"

"2" "1"

"x=\dfrac{2\sdot 0.01mol}{1}=0.02mol;"

"\nu(HCl)=0.02mol;"

"b) M(CaCl_2)=111g\/mol;"

"0.75g" "xg"

"2HCl+Ca(OH)_2 \to CaCl_2+2H_2O;"

"74g\/mol" "111g\/mol"

"x= \dfrac{0.75g\sdot111g/mol}{74g/mol}=1.125g;"

"m(CaCl_2)=1.125g;"

Answer: "a) \nu(HCl)=0.02mol;"

"b)m(CaCl_2)=1.125g."